1. It is known that 20% of all persons given a certain "medication" get very drowsy within 2.5 minutes. Find the probabilities that among 15
persons given the medications:
a. at most two will get very drowsy within 2 minutes.
b. at least five will get very drowsy within 2 minutes.

Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

at most 2 includes 0, 1, and 2
p(0)=15C0*0.8^15=0.0352
p(1)=15C1*0.2^1*0.8^14=0.1319
p(2)=15C2*0.2^2*0.8^13=0.2309
that total is 0.3980
Also binomcdf(15,0.2,2)
-at least 5 can do with 1-bionomcdf(15,0.2,4)=0.1642
OR do p(3)=15C3*0.2^3*0.8^12 and p(4)=15C4*0.2^4*0.8^11, sum those to the three above and subtract from 1.

Answer by ikleyn(52776)

.

As your problem is worded, printed and presented, it is DEFECTIVE.

The "given" part contains data 2.5 minute.

The questions asks about 2 minutes.


Worded this way, the "problem" is only good to throw it to the garbage bin.